OpenStudy (anonymous):

Describe an event that has probability of 0 and an event that has probability of 1.

OpenStudy (anonymous):

consider set of natural numbers. prob of selecting a number which is -ve from set is zero. prob of a prime number present in set is one

OpenStudy (anonymous):

if you roll a dice 1.whats the probability of getting a number above 6? its 0/6=0 2.whats the probability of getting a number between 0-7? its 6/6 = 1

OpenStudy (anonymous):

i like mathpro's but in the 2nd 1 write "integer" instead of "number"

OpenStudy (anonymous):

1.whats the probability of getting a number above 6? its 0/6=0 Is it the same as the probability of the dice landing on a corner?